A series RLC circuit has a capacitive reactance of 16 Ω and an inductive reactance of 4.0 Ω at 30 Hz. What frequency would result in resonance for this circuit?
Answer #1

X_{C}=\frac{1}{2\pi fC}=>C=\frac{1}{2\pi *16*30}=3,32x10^{-4}F

X_{L}=2\pi fL=>L=\frac{4}{2\pi *30}=0,021H

Resonance occurs when

X_{L}=X_{C} => 2\pi fL=\frac{1}{2\pi fC} => f^{2}=\frac{1}{4\pi ^{2}LC}

f=\sqrt{\frac{1}{4\pi ^{2}LC}}=\sqrt{\frac{1}{4\pi ^{2}(0,021)(3,32x10^{-4})}}=60Hz